Given the following model:
Consumption: C = 500 + 0.5Yd
Investment: I = 250
Government Expenditure: G = 100
Proportional Tax Rate: t = 0.1
Imports: M = 0.25Y
Exports: X = 50
Yd: Disposal Income
Y: Real GDP
(Note: There is no lump-sum tax)
a) If the current level of output is 1000, what is the level of actual investment? (2 marks)
b) Calculate the equilibrium real GDP (2 marks)
c) Given your answers to part (a) and (b), explain how the economy will adjust to short-
run equilibrium (3 marks)
d) Calculate the disposal income (Yd) in equilibrium (1 mark)
e) Calculate the multiplier (1 mark)
1
Expert's answer
2020-09-14T10:48:46-0400

a)

1000=500+Yd+I+100+50-0.25*1000

1000=500+0.5(1000-0.1*1000)+I+100+50-200

1000=500+450+I+100-150

I=100

b)Y=500+Yd+250+100+50-0.25Y

Y=900+0.5(Y-0.1Y)-0.25Y

Y-0.2Y=900

Y=4500

c)

In the short term, the economy is growing, which is a positive trend, but prices are likely to rise. Therefore, a restraining fiscal and monetary policy is necessary

d) Yd=Y-T=4500-0.1*4500=4050

e)"m=\frac{4500-1000}{250-100}=23.33"